OpenStudy (anonymous):

What is the exact value (using special triangles) of sin 4π/3 ?

OpenStudy (anonymous):

Do you know the exact value for sin(60)

OpenStudy (anonymous):

that is sin(pi/3)

OpenStudy (anonymous):

The associated acute angle of 4pi/3 is -pi/3. Hence \[\sin(4\pi/3)=\sin(-\pi/3)=-\sin(\pi/3)=-\frac{\sqrt 3}{2}\]

OpenStudy (anonymous):

Thankyou so much for your help! I have another question like this and I think I kind of get it